OpenStudy (anonymous):

Solve the system of equations and choose the correct ordered pair. 6x + 5y = -1 -7x + 15y = -103

OpenStudy (anonymous):

multiply first equation with 3 will get 18x+15y=-3 substract 18x+15y = -3 _ -7x +15y = -103 ________________________ 25x+ 0 = 100 __________________________ 25x=100 x=4 substitute x= 4 to get y 6x + 5y = -1 24+5y=-1 5y=-25 y=-5 answer: (4, -5)